# MRAMG-Bench: A BeyondText Benchmark for Multimodal Retrieval-Augmented Multimodal Generation

### v1

@@ 4.4. Data Statistics
The MRAMG-Bench, summarized in Table 2, comprises 4,800 QA pairs across three distinct domains: Web Data, Academic Paper Data, and Lifestyle Data. It integrates six datasets—MRAMG-Wit, MRAMG-Wiki, MRAMG-Web, MRAMG-Arxiv, MRAMG-Recipe, and MRAMG-Manual—containing 4,346 documents and 14,190 images, with tasks categorized into three difficulty levels. Datasets like MRAMG-Recipe and MRAMG-Manual feature longer texts and higher image density, posing greater challenges compared to others. The benchmark’s query distribution captures varying levels of complexity: Single-image QA primarily involves direct image-text associations. Two-image QA requires comparative analysis or step-wise reasoning. Three+ image QA focuses on complex multi-step processes. Text-only QA emphasizes pure textual reasoning and serves to evaluate the model’s ability to exclude irrelevant images when they are not required.
@@
DATASETS MRAMG-Wit MRAMG-Wiki MRAMG-Web MRAMG-Arxiv MRAMG-Recipe MRAMG-Manual Overall
@@
5. A General Framework for MRAMG
In this section, we present our generation framework, which consists of two stages: (1) the retrieval of relevant documents, and (2) the generation of multimodal answers based on the retrieved information, utilizing a foundational generative model. In the second stage, we propose three distinct strategies for answer generation: (a) LLM-based method: directly generating multimodal answers using LLMs, (b) MLLM-based method: directly generating multimodal answers using MLLMs, and (c) rule-based method: first obtaining a pure text answer from generative large models and then applying our matching-based algorithm to insert appropriate images into the text answer, resulting in multimodal answers. All the prompts we used can be found at: https://huggingface.co/MRAMG.
@@ 5.2.1. LLM-Based Method
We first consider how to generate multimodal answers using LLMs. Since LLMs cannot directly process image formats such as images for understanding, we choose an alternative method. In fact, we notice that the captions and surrounding context of images often offer detailed descriptions and supplementary information, which enhance the understanding of the images in the context of the query. Thus, we leverage the contextual information of images within the text context, as well as the captions of the images themselves, as textual substitute for the images when inputting into the language model. Specifically, we insert image placeholders $p_{1},p_{2},\dots,p_{n}$ at the corresponding positions of the retrieved text information $\mathcal{T}_{q}^{*}$, resulting in a new context $\mathcal{C}_{q}^{*}=\{t_{1},t_{2},\dots,t_{l_{1}},p_{1},t_{l_{1}+1},\dots,t_{l_{j}},p_{j},t_{l_{j}+1},\dots\}$. We consider that the context $t_{l_{j}}$ and $t_{l_{j}+1}$ around the placeholder $p_{j}$ for the $j$-th image $i_{j}$ provide a suitable textual description for that image. Then, we generate the multimodal answer as $\mathcal{A}=\mathcal{M}(\mathcal{P}_{llm},q,\mathcal{C}_{q}^{*})$, where $\mathcal{P}_{llm}$ is the prompt, and the images in the output are represented as placeholders.
@@ 5.2.2. MLLM-Based Method
For multimodal large language models (MLLMs), we directly input the textual information $\mathcal{T}_{q}^{*}$ and the image information $\mathcal{I}_{q}^{*}$. However, current open-source or closed-source MLLMs typically impose limits on the number of input images. Therefore, we are unable to use all the images in $\mathcal{I}_{q}^{*}$ as input. To address this, we propose leveraging the CLIP-based multimodal embedding model (Koukounas et al., 2024) to compute the embedding similarity between each image and the query, and then selecting the top-N images based on the similarity for input, denoted as $\operatorname{TopN}(q,\mathcal{I}_{q}^{*})$. Formally, for a given MLLM $\mathcal{M}$, we generate the multimodal answer as $\mathcal{A}=\mathcal{M}(\mathcal{P}_{mllm},\mathcal{T}_{q}^{*},\operatorname{TopN}(q,\mathcal{I}_{q}^{*}))$, where $\mathcal{P}_{mllm}$ is the prompt, and the images in the output are also represented as placeholders.
@@ 5.2.3. Rule-Based Method
Here, we consider an alternative strategy for generating multimodal answers. Specifically, we first allow the generative model to produce a pure textual answer based on the provided context. Then, we employ a matching-based insertion algorithm to select appropriate images from the retrieved set and insert them at suitable positions within the textual answer, thereby forming the multimodal answer.
@@
Step 2: Construct Bipartite Graph. We then consider constructing the following bipartite graph $\mathcal{B}=(\mathcal{S},\mathcal{I},\mathcal{E})$, where $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$ represents the set of representative vertices corresponding to the sentences in all pure text answers, and =$\mathcal{I}=\{i_{1},i_{2},\dots,i_{n}\}$ represents the set of representative vertices corresponding to each retrieved image. Here, we construct the edge set $\mathcal{E}$ by calculating the string similarity $\alpha_{k,l}$ (using BLEU method (Papineni et al., 2002)) and semantic similarity $\beta_{k,l}$ (using BGE-M3 model (Chen et al., 2024b)) for each sentence-image pair $(s_{k},i_{l})$. Specifically, to filter out irrelevant images, we define the edge set as =$\mathcal{E}=\{(k,l)\mid\alpha_{k,l}>\alpha^{*}\text{ or }\beta_{k,l}>\beta^{*}\}$, where $\alpha^{*},\beta^{*}$ are two hyper-parameters as thresholds, and each egde $(k,l)$ has the weight $w_{k,l}=\lambda\alpha_{k,l}+(1-\lambda)\beta_{k,l}$, where $\lambda\in[0,1]$ is a hyper-parameter.
@@
For the first constraint, we can solve for the optimal solution to problem (1) as $M^{*}=\left\{(k^{*},l)\mid k^{*}=\operatornamewithlimits{argmax}_{(k,l)\in\mathcal{E}}w_{k,l},\forall l\in[n]\right\}$. For the second constraint, $M$ forms a matching of the bipartite graph $\mathcal{B}$. We use the standard Edmonds’ Blossom algorithm (Edmonds, 1965) to solve the weighted maximum bipartite graph matching problem in this case. Finally, we insert the images into the output in the form of placeholders. We test the performance of the two matching-based insertion algorithms on the web and manual datasets, with results shown in Figure 3. Overall, the weighted bipartite graph matching algorithm, which restricts each sentence to have at most one image inserted after it, performed better. Therefore, we adopted this method for subsequent large-scale experimental evaluations.
@@ 6. Experiments
Table 3. Retrieval performance of the BGE-M3 model on the MRAMG-Bench.
@@
Framework Model Web Data Academic Paper Data Lifestyle Data
@@ 6.1.1. Retrieve Evaluation
  • Context Recall (Es et al., 2023) uses LLMs to evaluate whether the retrieved documents contain all the relevant information required for answer generation.
@@ 6.1.2. Generation Evaluation
To evaluate performance of multimodal answers, we consider the following metrics [^3] See detailed metrics at: https://huggingface.co/MRAMG ., which can be divided into two categories: statistical-based metrics (first six metrics) and LLM-based metrics (last four metrics).
@@
  • Image Ordering Score evaluates whether the order of images inserted into the multimodal answer matches the order of images in the ground truth  [^4] Here, we apply the ordering score only to lifestyle data (MRAMG-Recipe and MRAMG-Manual), as instances in these datasets typically contain a large number of images, and the order of the images is crucial.. Specifically, we compute the weighted edit distance between the two image sequences.
@@
  • Image Relevance measures evaluate the relevance of the inserted image to the query-answer pair, specifically assessing whether the content described by the image is meaningfully related to the content of the QA (Zhu et al., 2024a; Ma et al., 2024).
  • Image Effectiveness measures evaluate the effectiveness of the images inserted into the multimodal answer, assessing whether the images align with the QA content and contribute to the understanding of the answer (Zhu et al., 2024a; Ma et al., 2024).
@@
6.2. Experimental Baselines and Settings
To further validate our dataset and generation framework, we consider testing the generation performance of multimodal answers across a wide range of LLMs and MLLMs.
Model Baselines
We evaluate 4 closed-source models and 7 open-source models. Specifically, we select 4 popular closed-source multimodal large models: GPT-4o (OpenAI, 2024b), GPT-4o-mini (OpenAI, 2024a), Claude-3.5-Sonnet (Anthropic, 2024), Gemini-1.5-Pro (Team et al., 2024). For the open-source models, we choose 4 multimodal large models: Qwen2-VL-7B-Instruct (Wang et al., 2024a), Qwen2-VL-72B-Instruct (Wang et al., 2024a), InternVL-2.5-8B (Chen et al., 2024a), InternVL-2.5-78B (Chen et al., 2024a) and 3 text-based large models: DeepSeek-V3 (Liu et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024) and Llama-3.3-70B-Instruct (Dubey et al., 2024). Among these, for the four most widely used closed-source multimodal models, we test all the llm/mllm/rule-based strategies in our framework. For the four other multimodal models, we test the mllm/rule-based strategies. For the remaining three text-based models, we test the llm/rule-based strategies.
Experiment Details
@@ 6.3.1. Retrieval Performance
As shown in Table 3, we observe that both image recall and context recall perform well for Web Data , with MRAMG-Web achieving particularly high scores of 0.99 and 0.95, respectively. In contrast, both metrics progressively decline for the three other datasets , which can be attributed to their longer text passages and higher image counts. Nevertheless, overall retrieval effectiveness remains strong.
@@ 6.3.2. Generation Performance
In this section, we evaluate generation performance, with results presented for three domains (web/academic paper/lifestyle). Due to space limitations, full results are available at: https://huggingface.co/MRAMG.
@@
Generation Performance Across Different Datasets. From the results in Table 4, it is evident that the overall generation performance significantly decreases as the dataset complexity increases, which aligns with our expectations regarding the varying difficulty of different datasets. Notably, the performance of rule-based methods on simpler datasets is basically suboptimal, which can be attributed to data characteristics, particularly related to the Avg Images Per Doc metric. According to Table 2, this metric is 1 for all the Web Data datasets , whereas in the other datasets, it is substantially higher. With fewer images per document, the rule-based illustration algorithm struggles to effectively associate images with sentences, as multiple sentences often correspond to a single image. This results in ambiguous correlations and complicates threshold selection, leading to a higher error rate. In contrast, as the number of images increases, the gap in correlation between rule-based methods and other approaches becomes more pronounced. Under these conditions, rule-based matching performs better due to improved alignment with sentence-image relationships.
@@
Generation Performance Across Different Models. As shown in Table 4, advanced models such as Gemini, Claude, GPT-4o, and Deepseek-V3 consistently outperform smaller open-source models ($\sim 7$B parameters) across all domains and methods. These smaller models exhibit subpar performance across different methods and dataset domains, even when utilizing rule-based generation techniques. In contrast, larger open-source models ($\sim 70$B parameters) significantly reduce the performance gap with closed-source models, achieving results within approximately a tenth of the margin on simpler datasets, such as Web Data. For example, InternVL-2.5-78B and Llama-3.3-70B-Instruct attain Image Precision scores of 75.5 and 74.26, respectively, with average scores surpassing 70. On more challenging datasets, however, the performance gap becomes more pronounced, highlighting the limitations of open-source models in handling complex MRAMG tasks. Specifically, when employing MLLM-based methods, Qwen2-VL-72B-Instruct achieves an Image Recall score of only 26.25, which is significantly lower than Gemini’s 57.4. This result underscores the challenges faced by even 70B-scale models in accurately identifying images in complex scenarios. Nevertheless, smaller open-source models remain a cost-effective solution for simpler applications with limited computational resources. In particular, the order metric for the Lifestyle domain demonstrates poor performance across all models and methods, with none achieving a passing score. Notably, GPT-4o performs best under both the Rule-based and MLLM-based methods, scoring 43.54 and 32.47, respectively, while the Gemini model attains the highest score of 54.62 under the LLM-based method. However, even these scores fall short of the passing threshold, underscoring the ambiguity in LLM and MLLM models regarding the concept of image insertion order, which remains an unresolved challenge.
@@
LLM-based methods: By integrating contextual information surrounding images into the generation process, this method achieves natural and precise image insertion, underscoring the critical role of context in ensuring insertion accuracy.
MLLM-based methods: While effective on simpler datasets such as Web Data, their performance deteriorates on more challenging datasets due to the increased difficulty of distinguishing visually similar images, revealing the current limitations of model capabilities. In this category, the Gemini model achieves the highest average scores on the Web and Lifestyle Data, scoring 85.51 and 61.58, respectively, while Claude outperforms others on the Academic Paper Data with a score of 76.87.
Rule-based methods: Although these methods exhibit significantly lower performance compared to model-based approaches on simpler datasets, the performance gap diminishes as dataset complexity increases. In the Lifestyle Data domain, rule-based methods even outperform certain model-based approaches. For instance, GPT-4o achieves an average score of 65.68 using rule-based methods, surpassing its MLLM-based score of 61.01. Notably, the Deepseek-V3 model demonstrates outstanding performance on Web and Academic Paper Data, achieving top scores across almost all metrics. This suggests that for simpler datasets, rule-based methods align well with the Deepseek-V3 model, indicating the model’s tendency to leverage such approaches effectively. Despite their limitations, rule-based methods offer several key advantages:
@@
Overall, rule-based methods provide a viable and efficient alternative, particularly in resource-constrained or stability-critical scenarios. Furthermore, the performance of LLM-based methods is generally on par with or even superior to rule-based methods, showcasing the strong in-context reasoning capabilities of modern large models.

### v2

@@ 4.4. Data Statistics
Statistic MRAMG-Wit MRAMG-Wiki MRAMG-Web MRAMG-Arxiv MRAMG-Recipe MRAMG-Manual Overall
@@
The MRAMG-Bench, summarized in Table 2, comprises 4,800 QA pairs across three distinct domains: Web, Academia, and Lifestyle. It integrates six datasets—MRAMG-Wit, MRAMG-Wiki, MRAMG-Web, MRAMG-Arxiv, MRAMG-Recipe, and MRAMG-Manual—containing 4,346 documents and 14,190 images, with tasks categorized into three difficulty levels. Datasets like MRAMG-Recipe and MRAMG-Manual feature longer texts and higher image density, posing greater challenges compared to others. The benchmark’s query distribution captures varying levels of complexity: Single-image QA primarily involves direct image-text associations. Two-image QA requires comparative analysis or step-wise reasoning. Three+ image QA focuses on complex multi-step processes. Text-only QA emphasizes pure textual reasoning and serves to evaluate the model’s ability to exclude irrelevant images when they are not required.
5. Generation Framework for MRAMG Task
In this section, we present our generation framework, which consists of two stages: (1) the retrieval of relevant documents, and (2) the generation of multimodal answers based on the retrieved information, utilizing a foundational generative model. In the second stage, we propose three distinct strategies for answer generation: (a) LLM-based method: directly generating multimodal answers using LLMs, (b) MLLM-based method: directly generating multimodal answers using MLLMs, and (c) rule-based method: first obtaining a pure text answer from generative large models and then applying our matching-based algorithm to insert appropriate images into the text answer, resulting in multimodal answers. All the prompts we used can be found at https://github.com/MRAMG-Bench/MRAMG.
@@ 5.2.1. LLM-Based Method
We first consider how to generate multimodal answers using LLMs. Since LLMs cannot directly process image formats such as images for understanding, we choose an alternative method. In fact, we notice that the captions and surrounding context of images often offer detailed descriptions and supplementary information, which enhance the understanding of the images in the context of the query. Thus, we leverage the contextual information of images within the text context, as well as the captions of the images themselves, as textual substitute for the images when inputting into the language model. Specifically, we insert image placeholders $p_{1},p_{2},\dots,p_{n}$ at the corresponding positions of the retrieved text information $\mathcal{T}_{q}^{*}$, resulting in a new context $\mathcal{C}_{q}^{*}=\{t_{1},t_{2},\dots,t_{l_{1}},p_{1},t_{l_{1}+1},\dots,t_{l_{j}},p_{j},t_{l_{j}+1},\dots\}$. We consider that the context $t_{l_{j}}$ and $t_{l_{j}+1}$ around the placeholder $p_{j}$ for the $j$-th image $i_{j}$ provide a suitable textual description for that image. Then, for a given LLM $\mathcal{M}$, we generate the multimodal answer as $\mathcal{A}=\mathcal{M}(\mathcal{P}_{llm},q,\mathcal{C}_{q}^{*})$, where $\mathcal{P}_{llm}$ is the prompt, and the images in the output are represented as placeholders.
@@ 5.2.2. MLLM-Based Method
For MLLMs, we can directly input the textual information $\mathcal{T}_{q}^{*}$ and the image information $\mathcal{I}_{q}^{*}$. However, current open-source or closed-source MLLMs typically impose limits on the number of input images. Therefore, we are unable to use all the images in $\mathcal{I}_{q}^{*}$ as input. To address this, we propose leveraging the CLIP-based multimodal embedding model (Koukounas et al., 2024) to compute the embedding similarity between each image and the query, and then selecting the top-N images based on the similarity for input, denoted as $\operatorname{TopN}(q,\mathcal{I}_{q}^{*})$. Formally, for a given MLLM $\mathcal{M}$, we generate the multimodal answer as $\mathcal{A}=\mathcal{M}(\mathcal{P}_{mllm},\mathcal{T}_{q}^{*},\operatorname{TopN}(q,\mathcal{I}_{q}^{*}))$, where $\mathcal{P}_{mllm}$ is the prompt, and the images in the output are also represented as placeholders.
@@ 5.2.3. Rule-Based Method
Here, we consider an alternative strategy for generating multimodal answers. Specifically, we first leverage the generative model to generate a pure textual answer based on the provided context. Then, we employ a matching-based insertion algorithm to select appropriate images from the retrieved set and insert them at suitable positions within the answer, thereby forming the multimodal answer. It is carried out in three steps:
@@
Step 2: Construct Bipartite Graph. We then consider constructing the following bipartite graph $\mathcal{B}=(\mathcal{S},\mathcal{I},\mathcal{E})$, where $\mathcal{S}=\{s_{1},s_{2},\dots,s_{m}\}$ represents the set of representative vertices corresponding to the sentences in all pure text answers, and =$\mathcal{I}=\{i_{1},i_{2},\dots,i_{n}\}$ represents the set of representative vertices corresponding to retrieved images. Here, we construct the edge set $\mathcal{E}$ by calculating the string similarity $\alpha_{k,l}$ (using BLEU method (Papineni et al., 2002)) and semantic similarity $\beta_{k,l}$ (using BGE-M3 model (Chen et al., 2024b)) for each sentence-image pair $(s_{k},i_{l})$. Specifically, to filter out irrelevant images, we define the edge set as =$\mathcal{E}=\{(k,l)\mid\alpha_{k,l}>\alpha^{*}\text{ or }\beta_{k,l}>\beta^{*}\}$, where $\alpha^{*},\beta^{*}$ are two hyper-parameters as thresholds, and each egde $(k,l)$ has the weight $w_{k,l}=\lambda\alpha_{k,l}+(1-\lambda)\beta_{k,l}$, where $\lambda\in[0,1]$ is a hyper-parameter.
@@
For the first constraint, we can solve for the optimal solution to problem (1) as (Max Similarity Algorithm):
(2) $M^{*}=\left\{(k^{*},l)\mid k^{*}=\operatornamewithlimits{argmax}_{(k,l)\in\mathcal{E}}w_{k,l},\forall l\in[n]\right\}.$
For the second constraint, $M$ forms a matching of the bipartite graph $\mathcal{B}$. We use the standard Edmonds’ Blossom algorithm (Edmonds, 1965) to solve the weighted maximum bipartite graph matching problem in this case (Graph Matching Algorithm).
Finally, we insert the images into the output in the form of placeholders. We test the performances of the two matching-based insertion algorithms on the web and manual datasets, with results shown in Figure 3. Overall, the Graph Matching Algorithm, which restricts each sentence to have at most one image inserted after it, performs better. Therefore, we adopt this method for subsequent large-scale experimental evaluations.
@@ 6. Experiments
6.1. Experimental Baselines and Settings
Baselines
We evaluate 4 closed-source models and 7 open-source models. Specifically, we select 4 popular closed-source multimodal large models: GPT-4o (OpenAI, 2024b), GPT-4o-mini (OpenAI, 2024a), Claude-3.5-Sonnet (Anthropic, 2024), Gemini-1.5-Pro (Team et al., 2024). For the open-source models, we choose 4 multimodal large models: Qwen2-VL-7B-Instruct (Wang et al., 2024a), Qwen2-VL-72B-Instruct (Wang et al., 2024a), InternVL-2.5-8B (Chen et al., 2024a), InternVL-2.5-78B (Chen et al., 2024a) and 3 text-based large models: DeepSeek-V3 (Liu et al., 2024), Llama-3.1-8B-Instruct (Dubey et al., 2024) and Llama-3.3-70B-Instruct (Dubey et al., 2024). Among these, for the four most widely used closed-source multimodal models, we test all the LLM/MLLM/rule-based methods in our framework. For the four other multimodal models, we test the MLLM/rule-based methods. For the remaining three text-based models, we test the LLM/rule-based methods.
Experiment Details
Table 3. Retrieval performance of the BGE-M3 model on MRAMG-Bench.
@@ 6.1. Experimental Baselines and Settings
Framework Model Web Data Academic Data Lifestyle Data
@@ 6.2.1. Retrieval Evaluation
  • Context Recall (Es et al., 2023) uses LLMs to evaluate whether the retrieved documents contain all the relevant textual information required for answer generation.
@@ 6.2.2. Generation Evaluation
To evaluate performance of multimodal answers, we consider the following metrics [^3] See detailed metrics at https://github.com/MRAMG-Bench/MRAMG ., which can be divided into two categories: statistical-based metrics (first six metrics) and LLM-based metrics (last four metrics).
@@
  • Image Ordering Score evaluates whether the order of images inserted into the multimodal answer matches the order of images in the ground truth  [^4] Here, we apply the ordering score only to Lifestyle Data , as instances in these datasets typically contain a large number of images, and the order of the images is crucial.. Specifically, we compute the weighted edit distance between the two image sequences.
@@
  • Image Relevance measures the relevance of the inserted image to the query-answer pair, specifically assessing whether the content described by the image is meaningfully related to the content of the QA (Zhu et al., 2024a; Ma et al., 2024).
  • Image Effectiveness measures the effectiveness of the images inserted into the multimodal answer, assessing whether the images align with the QA content and contribute to the understanding of the answer (Zhu et al., 2024a; Ma et al., 2024).
@@ 6.3.1. Retrieval Performance
As shown in Table 3, we observe that both visual recall and context recall perform well for Web Data , with MRAMG-Web achieving particularly high scores of 0.99 and 0.95, respectively. In contrast, both metrics progressively decline for the three other datasets , which can be attributed to their longer text passages and higher image counts. Nevertheless, overall retrieval effectiveness remains strong.
@@ 6.3.2. Generation Performance
In this section, we evaluate generation performance, with results presented for three domains (Web/ Academia/ Lifestyle). Due to space limitations, complete results are available at https://github.com/MRAMG-Bench/MRAMG.
@@
Generation Performance Across Different Datasets. From the results in Table 4, it is evident that the overall generation performance significantly decreases as the dataset complexity increases, which aligns with our expectations regarding the varying difficulty of different datasets. Notably, the performance of rule-based method on simpler datasets is basically suboptimal, which can be attributed to data characteristics, particularly related to the Avg Images Per Doc metric. According to Table 2, this metric is 1 for all the web datasets , whereas in the other datasets, it is substantially higher. With fewer images per document, the rule-based method struggles to effectively associate images with sentences, as multiple sentences often correspond to a single image. This results in ambiguous correlations and complicated threshold selection, leading to a higher error rate. In contrast, as the number of images increases, the gap in correlation between rule-based method and other methods becomes more pronounced. Under these conditions, rule-based performs better due to improved alignment with sentence-image relationships.
@@
Generation Performance Across Different Models. As shown in Table 4, advanced models such as Gemini, Claude, GPT-4o, and Deepseek-V3 consistently outperform smaller open-source models ($\sim 7$B parameters) across all domains and methods. These smaller models exhibit subpar performance across different methods and dataset domains, even when utilizing rule-based generation techniques. In contrast, larger open-source models ($\sim 70$B parameters) significantly reduce the performance gap with closed-source models, achieving results within approximately a tenth of the margin on simpler datasets, such as Web Data. For example, InternVL-2.5-78B and Llama-3.3-70B-Instruct attain Image Precision scores of 75.5 and 74.26, respectively, with average scores surpassing 70.
On more challenging datasets, however, the performance gap becomes more pronounced, highlighting the limitations of open-source models in handling complex MRAMG tasks. Specifically, when employing MLLM-based method, Qwen2-VL-72B-Instruct achieves an Image Recall score of only 26.25 on Lifestyle Data, which is significantly lower than Gemini’s 57.4. This result underscores the challenges faced by even 70B-scale models in accurately identifying images in complex scenarios. Nevertheless, smaller open-source models remain a cost-effective solution for simpler applications with limited computational resources.
In particular, the order metric for the lifestyle domain demonstrates poor performance across all models and methods, with none achieving a passing score. Notably, GPT-4o performs best under both the rule-based and MLLM-based methods, scoring 43.54 and 32.47, respectively, while the Gemini model attains the highest score of 54.62 under the LLM-based method. However, even these scores fall short of the passing threshold, underscoring the ambiguity in LLM and MLLM models regarding the concept of image insertion order, which remains an unresolved challenge.
@@
LLM-Based Method: By integrating contextual information surrounding images into the generation process, this method achieves natural and precise image insertion, underscoring the critical role of context in ensuring insertion accuracy.
MLLM-Based Method: While effective on simpler datasets such as Web Data, their performance deteriorates on more challenging datasets due to the increased difficulty of distinguishing visually similar images, revealing the current limitations of model capabilities. In this category, the Gemini model achieves the highest average scores on Web and Lifestyle Data, scoring 85.51 and 61.58, respectively, while Claude outperforms others on Academic Data with a score of 76.87.
Rule-Based Method: Although these methods exhibit significantly lower performance compared to model-based approaches on simpler datasets, the performance gap diminishes as dataset complexity increases. In the lifestyle domain, rule-based methods even outperform certain model-based approaches. For instance, GPT-4o achieves an average score of 65.68 using rule-based method, surpassing its MLLM-based score of 61.01. Notably, the Deepseek-V3 model demonstrates outstanding performance on Web Data and Academic Data, achieving top scores across almost all metrics. This suggests that for simpler datasets, rule-based method aligns well with the Deepseek-V3 model, indicating the model’s tendency to leverage such approaches effectively. Despite their limitations, rule-based methods offer several key advantages:
@@
Overall, rule-based methods provide a viable and efficient alternative, particularly in resource-constrained or stability-critical scenarios. Furthermore, the performance of LLM-based methods is generally on par with or even superior to rule-based methods, showcasing the strong in-context reasoning capabilities of large models.